find its equation: having the same y-intercept as 2x+5y=-25 and twice the slope of 9x-3y=4
Answer by mananth(16946)

find its equation: having the same y-intercept as 2x+5y=-25 and twice the slope of 9x-3y=4

find its equation: having the same y-intercept as 2x+5y=-25
5y=-2x-25
y= (-2/5)x -5
y intercept =-5

and twice the slope of 9x-3y=4
3y=9x-4
y= 3y-(4/3)
m=3
2m =6
The required equation
y = 6x-5